[Halld] Commissioning and shift updates

Yi Qiang yqiang at jlab.org
Thu Nov 13 11:01:50 EST 2014


Dear Collaborators,

The Cryo group is still working on solving the 2k Cold Box issue at this 
moment so there will be no beam until at least 2 PM today. Therefore, 
all shifts to 2 PM today remain stand-by and I will keep you updated 
with the latest status.

After the electron beam is restored, our first priority is to 
re-establish the photon beam in the hall and calibrate the active 
collimator. More detailed plan will be posted on the HDLOG logbook shortly.

Best,

Yi Qiang, Hall D Run Coordinator.

RC Cell: 757-383-5542


More information about the Halld mailing list